OpenStudy (anonymous):

Can someone solve this problem for me please? Find the quadratic polynomial whose graph goes through the points (-1,1), (0,2), and (1,7)?

OpenStudy (anonymous):

A quadratic polynomial has the form \[y = ax^{2}+bx+c\] Since you have 3 solution points, substitute each of them into the polynomial above: For (0,2): \[2 = a(0)^{2} + b(0) + c = c\] For (-1,1): \[1 = a(-1)^{2} + b(-1) + c = a - b + c = a - b + 2\] For (1,7): \[7 = a(1)^{2} + b(1) + c = a + b + c = a + b + 2\] Now you have two equations: \[1 = a - b + 2\] and \[7 = a + b + 2\] Simplify: \[a - b = -1\] and \[a + b = 5\] Solve the system of equations by substitution: \[a = b - 1\] \[a + b = (b - 1) + b = 2b - 1 = 5\] \[b = 3\] and \[a = 2\] Rewrite the quadratic polynomial with the constants: \[y = 2x^{2} + 3x + 2\]

OpenStudy (anonymous):

Thank you so much